# Recurrent convolutional neural networks for non-adiabatic dynamics of quantum-classical systems

### v1

@@ III.2 Implementation and training details
In accordance with [49], we utilize channel-wise dropout instead of element-wise dropout. Traditional element-wise dropout, when placed between two layers of a neural network, helps to regularize by setting a random proportion of incoming values to zero (hence “dropout”) during training. This forces the next layer to use a different subset of the incoming information each time by reducing reliance on only specific parts of the information. However, this random dropout of all values can introduce noise to convolution layers by interfering with local relationships, thereby reducing overall effectiveness. Channel-wise dropout addresses this by dropping out entire channels in a convolution layer, thereby fully preserving important locality information in the remaining channels.
IV Results
FIG. 2 shows the comparison of PARC simulation and ground truth for the time evolution of a $L=8$ system subject to a shallow quench from $g_{i}=0.5$ to $g_{f}=0.8$ at time $t=0$. The ground truth was obtained from integration with the 4th-order Runge-Kutta (RK4) method. A time-step $\Delta t=0.08$ was used both in the RK4 integration and the PARC prediction; the simulation from $t=0$ to $t=32$ thus corresponds to 400 iterations of predictions from the model. As discussed above, the system is initially in a CDW state where the on-site electron number $n_{i}=\rho_{ii}$ exhibits an alternating $\overline{n}+\delta n$ and $\overline{n}-\delta n$ pattern along the chain with the period of one lattice constant. This is accompanied by a ±$\pm Q_{0}$ modulation due to the finite electron-phonon coupling $g_{i}$. After the quench, the relaxation of the system towards the shifted ground state induces a coherent oscillation. Importantly, during this coherent regime, the system remains in a perfect CDW state with a time-varying modulation amplitude. Our benchmark shows that the PARC predictions accurately captures the coherent oscillation dynamics in all three $Q$, $P$, and $\rho$ sectors.
Next we consider the deep quench scenario, where an initial state of decoupled electrons and phonons (corresponding to $g_{i}=0$) is perturbed by a switch-on of the electron-phonon coupling. The comparison of the ground truth and the PARC prediction is shown in FIG. 3. To demonstrate the generalization capability of the PARC model, an arbitrary nonequilibrium state for both phonon and electron subsystems is used as the initial state. Contrary to a persistent long-range CDW order in the shallow quench case, the dynamical evolution is characterized by the presence of strong spatial fluctuations. The spatial inhomogeneity also manifests itself in the localization of electrons, giving rise to relatively small off-diagonal matrix elements in the electron density matrix $\rho$. As the figure shows, the rather inhomogeneous dynamical evolution is also nicely captured by the PARC model. It is also worth noting that this benchmark was done on an initial state which was not among the training dataset. This thus demonstrates the transferability of the PARC model.
Next we present systematic studies on model prediction accuracy with respect to prediction window, prediction step size, training dataset size, and complexity of NN models. To quantify the prediction accuracy, we define the RMSE of each of $\rho$, $Q$, and $P$ on next-step prediction as follows:
$\text{RMSE}_{Q}$ $=\sqrt{\frac{\left[Q(t+\Delta t)-Q_{\text{pred}}(t+\Delta t)\right]^{2}}{L}}$
$\text{RMSE}_{P}$ $=\sqrt{\frac{\left[P(t+\Delta t)-P_{\text{pred}}(t+\Delta t)\right]^{2}}{L}}$
$\text{RMSE}_{\rho}$ $=\sqrt{\frac{|\rho(t+\Delta t)-\rho_{\text{pred}}(t+\Delta t)|^{2}}{L^{2}}}$
Where $L$ denotes the system size (the square size of $\rho$ and length of $Q$ and $P$) and ||$\cdot$|| ⋅ | denotes element-wise magnitude. Then the total RMSE of next-step prediction is defined as:
$\text{RMSE}=\text{RMSE}_{\rho}+\text{RMSE}_{Q}+\text{RMSE}_{P}$
FIG. 4 shows the increase in prediction RMSE with respect to the number of steps predicted for. We can see that initial error accumulation appears to be relatively linear, but this trends exponential as the error continues to compound for greater number of steps. The magnitude of error for all step counts is around $2$ times lower for model trained on shallow-quench data in FIG. 4(a) compared to the model trained on deep-quench data in FIG. 4(b).
Figure 4: RMSE versus prediction window for (a) shallow quench $g=0.5$ to $0.8$ and (b) a deep quench with $g=0$ to 2.0.
Figure 5: Total RMSE versus the prediction step size $\Delta t$ for the PARC model trained on the deep quench data.
This can be explained as the model is more capable of memorizing and overfitting to the single trajectory of the small quench scenario, whereas generalization is vital for the performance of the deep-quench model. As a result, the shallow-quench trained PARC model, having effectively memorized the single trajectory data, performs better on the same data than the deep-quench trained model on a test set it was not trained on. However, being less generalized itself, the single-trajectory PARC model is also less capable of handling out-of-distribution situations, which in this case takes the form of its predictions as error accumulates with step size, resulting in increasingly unfamiliar inputs. As a result, we see that while the errors of the single-trajectory model may be lower for all step counts tested, the scaling of its error with respect to step count is more exponential than that of the deep-quench trained case, which can better handle error accumulation due to increased generalization on a variety of trajectories.
We next test the scaling of the RMSE of the PARC model trained on the deep-quench data with step sizes of $\Delta t=0.02,0.04,0.08,0.16$. FIG. 5 displays the average RMSE of a model trained on a data set with a certain step size on a corresponding test set. As expected, the error increases with step size. We can see that the scaling is somewhat greater than linear, which is expected as the degree of uncertainty compounds with larger step sizes.
Figure 6: Total RMSE vs size of training dataset measured by the number of deep-quench trajectories used in the training.
We also experiment with training the model on datasets containing $32$, $64$, $128$, $256$, $512$, $1024$, and $2048$ trajectories. All models were trained with the same amount of compute to enable an even comparison. As we can see in FIG. 6, all datasets except that with $32$ trajectories perform roughly evenly, indicating a sharp asymptote in the benefit of increased data diversity to generalization between datasets containing $32$ trajectories to those containing $64$. This shows that there is minimal practical gain to using datasets containing more than 64 trajectories, although there does not seem to be any harm in increased diversity either. The lack of harm from increased dataset size seems to imply that the model generalizes well, as it does not rely on memorization of the individual trajectories.
Finally, we examine the dependence of the prediction errors on the complexity of the PARC model. To this end, we scale the parameter count of the model by proportionally scaling the number of channels within each convolution layer of PARC. By default, PARC has $128$ channels in its convolution with the greatest number of channels. We test scaling which results in $32$, $64$, $128$, and $256$ channels at the convolution with greatest channel count. FIG. 7 shows the RMSE of such models on a test set against their total parameter count. As expected, the RMSE of the model decreases as the capacity of the model increases. This indicates that increasing the size of the model is an avenue for increased performance. However, we can also see from the figure, which has a logarithmic horizontal axis, that the benefit derived from increased size exhibits sublinear scaling and, as the total RMSE of prediction can only ever approach 00, exhibits diminishing returns as parameter count increases. Beyond a certain point it is likely ineffective or impractical to continue scaling the model due to the trade-off between performance and training/inference.
V Summary and Outlook
Figure 7: Total RMSE vs parameters of the CNN in the PARC model.
We have demonstrated the efficacy of a PARC-based recurrent CNN model to capture the time evolution of a one-dimensional semiclassical Holstein model under a quantum interaction quench. Within the Ehrenfest dynamics formalism, the non-adiabatic dynamics of the Holstein model is governed by the classical equation of motion for phonons, described by a set of simple harmonic oscillators and the von Neumann equation for the reduced density matrix of electrons. Leveraging the capability of CNNs to learn hierarchical local spatial correlations, our model is capable of generalizing to the integration of solutions to unseen initial conditions. The task of integration approximation is further augmented by the explicit differentiator-integrator structure of the PARC model, which integrates a physics-aware prior into the architecture. By utilizing a recurrent, next-step prediction target, our model is capable of predicting many steps of integration given any starting point.
Our benchmark studies on two quench scenarios have also demonstrated the accuracy of predicted trajectories to ground-truth. Because our model utilizes three separate CNNs for the density matrix, displacement vector, and momentum vector in the Holstein system, this indicates that each CNN has separately learned to appropriately capture the rules of integration for its corresponding equation in the system of PDEs from only the data. Additionally, we have explored the error accumulation with continued recurrent predictions and scaling properties of the step size of the data, size of the training set, and size of the model on the error of the prediction. Through these analyses we have been able to deduce effective model configurations and areas of diminishing returns.
A current area for improvement could be the long-term prediction accuracy of the model, and this could serve as a valuable subject for extension. For example, techniques used to increase the effective context window of RNNs could be applied, or even hybrid approaches such as integrating weak PINN constraints to the system.
[…]

### v3

@@ III.2 Neural network implementation
Second, we introduce channel-wise dropout [53] layers preceding the final linear layer of each residual block in the CNN. These layers randomly zero a certain proportion of channels during training, and act as a regularization technique which prevents the model from depending too heavily on any particular channel. Although the original ResNet-v2 structure does not employ any form of dropout, we find its addition to be critical in improving the model’s accuracy and stability. We utilize channel-wise dropout instead of traditional element-wise dropout - which works by zeroing a random proportion of all incoming values to zero - because the random dropout of all values can introduce noise to convolutional layers by interfering with local relationships, thereby reducing overall effectiveness. Channel-wise dropout addresses this by dropping out entire channels in a convolutional layer, thereby fully preserving important locality information in the remaining channels.
Third, the Tanh (hyperbolic tangent) activation function was used instead of the original ReLU (rectified linear unit) in ResNet-v2. This was due to observed beneficial effects of Tanh to model stability and accuracy. This may be a result of Tanh being both smooth and bounded while ReLU is neither; as a result, Tanh may provide a more suitable inductive bias for regression tasks which are both smooth and empirically bounded, such as approximating integration on the Holstein model.
Along with the aforementioned modifications, we also employ circular padding in our convolutional layers, as this naturally fits the periodic boundary conditions used for the simulations of the Holstein model in all state variables. Additionally, all inputs/outputs of the model contain data scaling coefficients. These are simple scalar constants set at the beginning of training based on training data statistics which normalize inputs and outputs such that the greatest absolute value in the training dataset should be $1$ [54]. Normalization of inputs is common practice which can help to ensure the different components of the input have the same scale and that the range of the input is located near the region of the activation function that is most expressive. Coefficients are set separately for $\rho$, $Q$, and $P$. For example, if the largest absolute value in $\rho$ within the training dataset is found to be $0.88$, the largest absolute value in $Q$ is $1.62$, and the largest absolute value in $P$ is $0.75$, then all inputs to the CNN would have the $\rho$ component scaled by $\frac{1}{0.88}$, the $Q$ component scaled by $\frac{1}{1.62}$, and the $P$ component scaled by $\frac{1}{0.75}$. These scaling coefficients were not shown in figure 2 to reduce complexity, however, more details and a diagram can be found in appendix A.
Figure 3: Snapshots of phonon displacements $Q_{i}$, momentum $P_{i}$, and electron density matrix $\rho_{ij}$ at various times after a shallow quench from $g_{i}=0.5$ to $g_{f}=0.8$ of the semi-classical Holstein model on a $L=16$ chain. The ground truth on the top row is obtained from the 4th-order Runge-Kutta integration of the governing dynamics Eqs. (3) and (6). The bottom row shows predictions from the trained simple CNN model. Only the real component of the complex-valued $\rho$ is shown. The steps are in units of prediction time steps, which for the shallow quench case is set to $\Delta t=0.64$ time units. The model is given the system state at prediction step 0 and predicts for the following $48$ prediction steps.
III.3 Shallow quench data generation
The training dataset was generated from direct numerical simulations of the semi-classical Holstein model. Specifically, the coupled Newton equation (3) for classical phonons and the von Neumann equation (6) for the electron density matrix were integrated using the standard fourth-order Runge-Kutta method. We pair the shallow quench scenario with the standard CNN to illustrate the relative simplicity of this regime, while the more complex deep quench scenario is modeled using the PARC-based approach, as discussed in Section IV.
In the shallow quench scenario, the system is initially prepared in the CDW ground state stabilized by a finite electron-phonon coupling $g_{i}>0$, which is then suddenly increased at $t=0$ to a final $g_{f}>g_{i}$ with a relatively small ratio $g_{f}/g_{i}\gtrsim 1$. This abrupt change induces transient coherent oscillations of the CDW order parameter, a nonequilibrium phenomenon observed in interaction quenches of various symmetry-breaking systems, including CDW, spin-density wave, and superconducting states. Importantly, the subsequent time evolution after a shallow quench is nearly deterministic. Consequently, this dataset provides an upper bound for the model’s ability to learn simple trajectories, serving as a benchmark for assessing its generalization performance.
To generate shallow quench data, a quench from $g_{i}=0.5$ to $g_{f}=0.8$ was used. The integration time step was set to $\delta t=0.01$ time units. Because the machine learning model can learn to predict a time step longer than the integration time step, a single prediction time step was saved every $64$ integration time steps; ie. $\Delta t=64\,\delta t=0.64$. This prediction time step length was chosen to reasonably balance efficiency with effectiveness; a longer prediction interval results in less compute when using the model to approximate the integration over a given time interval, however it may become difficult for the model to learn underlying dynamics if the prediction interval is too long. We present an analysis of the effects of the prediction interval on prediction error in appendix B. A total of $1201$ prediction time steps were saved for each trajectory, representing an initial state followed by $1200$ prediction time steps. Accordingly, this represented a domain spanning $1200\,\cdot\,0.64=768$ time units. We captured $64$ trajectories in total, each offset by one integration time step to provide uniform temporal sampling throughout the prediction interval of $\Delta t=0.64$.
III.4 Training details
Given the dataset in the form of exact system trajectories $\mathbf{u}(t)=[\rho(t),Q(t),P(t)]$, the loss function defined in Eq. (III.1) is used to obtain the optimized parameters $\bm{\theta}^{*}$ in the standard CNN model case. As can be seen in the loss function, the overall loss is the sum over multiple time steps. This is because we use multi-step prediction while training, wherein for some step value $N\geq 1$, the model recurrently generates predictions for prediction time steps $t+\Delta t,t+2\Delta t,\dots,t+N\Delta t$ when given the input at prediction time step $t$. The resulting predictions for the $N$ prediction time steps are then compared to the label values for the corresponding $N$ prediction time steps in the data for the calculation of the loss. We find that employing multi-step prediction greatly improves the model’s accuracy, likely because it forces the model to adapt to its own errors and better learn the long-term dynamics in the data.
The AdamW optimizer [55], which is a variant of the popular Adaptive Moment Estimation (Adam) algorithm [56], is used for gradient descent minimization of the loss function. AdamW is a simple modification of Adam which handles the weight decay - a technique to preventing overfitting by penalizing large weights in neural nets - correctly. Weight decay helps improve model generalization as large weight values can often indicate a model is ”relying” on specific, memorized features within training data instead of properly generalizing.
Alongside AdamW and weight decay, a few additional methods are used to stabilize and improve training. We use gradient clipping, a common technique which clips the gradient calculated during training to a maximum norm threshold to prevent massive and destabilizing updates to the model weights. This is an especially notable concern while training RNNs, as training with multi-step prediction can result in exploding gradients (gradients which accumulate to enormous values) [57].
During training, we add a small amount of gaussian noise to the input of the model. The motivation behind this is to improve the model’s ability to adapt to and correct deviations from the training data distribution, which is unavoidable as predictions errors compound during long recurrent prediction sequences. We find that the addition of this noise during training is highly beneficial to the model’s generalization abilities.
We employ curriculum learning [58] while training, which gradually increases the difficulty of the training process as the model learns. Specifically, we gradually increase the magnitude of the gaussian noise added to the input and the number of steps used in multi-step prediction with each stage of the curriculum. To compliment curriculum learning, we use a learning rate scheduler (which modifies the learning rate during training) which implements cosine annealing with warm restarts [59], and we align the restarts with the stages of our curriculum. Essentially, this learning rate scheduler maximizes the learning rate at the start of each stage in the curriculum and then slowly decreases the learning rate following a cosine curve, allowing the model to quickly learn at the start and then slow down updates as its weights converge. In addition, we further augment the learning rate scheduler by using a short linear learning rate warm-up at the beginning of each restart, which warms up the learning rate to its maximum value over a certain number of steps instead of at once. This is beneficial because Adam-family optimizers rely on estimates of the first and second moments of the gradient, and utilizing a warm-up reduces the updates the optimizer can make until it has learned better estimates for these values
Figure 4: Graphed ground truth (blue) vs predicted (red) $\Delta_{\rho}$ and $\Delta_{Q}$ for a shallow quench trajectory with system size $32$. The red prediction essentially perfectly covers the blue ground truth, which is under the red prediction in the visualization.
III.5 Benchmark
FIG. 3 illustrates both the ground-truth trajectories and the predictions generated by a standard CNN model trained on shallow-quench data for a system of size $16$. Owing to the highly deterministic and nearly periodic nature of the shallow-quench dynamics, we find that even a very compact CNN model - on the order of $\sim 5$K trainable parameters - suffices to fully capture the system’s evolution.
To present a more quantitative comparison, we first introduce the time-dependent order parameters of the system. First, the CDW order, or electron-number modulation, with a wave-vector $Q=\pi/a$ can be described by
$\Delta_{\rho}(t)=\frac{1}{L}\sum_{i}\rho_{ii}(t)\cos\left(\frac{\pi}{a}x_{i}\right).$ (13)
The electron number expectation value here is directly given by the diagonal elements of the density matrix: $\langle\hat{n}_{i}\rangle=\rho_{ii}(t)$. The ideal CDW order described by Eq. (2) corresponds to an order parameter of $\Delta=\delta n$. The staggered lattice distortion accompanying the charge modulation can be described by a similar order parameter
$\Delta_{Q}(t)=\frac{1}{L}\sum_{i}Q_{i}(t)\cos\left(\frac{\pi}{a}x_{i}\right).$ (14)
The time evolution of the two order parameters is shown in FIG. 4, comparing exact numerical simulations (ground truth) with CNN predictions. The Holstein system is initialized with the exact state and evolved forward for $1200$ prediction time steps. The CNN reproduces the periodic oscillations with essentially perfect fidelity, showing no visible deviation from the ground truth.
This striking accuracy reflects the relative simplicity of the shallow-quench regime: the deterministic dynamics can be effectively “memorized” by a small, standard CNN. In this case, the learning task requires neither elaborate architecture nor large model capacity, underscoring that the shallow quench serves primarily as a baseline demonstration rather than a demanding test of predictive power.
Figure 5: Schematic diagram of the recurrent structure for non-adiabatic dynamics of the semi-classical Holstein model. The framework is based on a single PARC-based CNN differentiator (diff)/integrator (int) pair for all three dynamical degrees of freedom: $Q$, $P$, and $\rho$.
IV PARC CNN and Deep Quench
Although a recurrent structure based on a single standard CNN suffices to reproduce the relatively deterministic dynamics of the Holstein model following a shallow quench, capturing more complex chaotic spatiotemporal evolution motivates the incorporation of additional physical constraints. To address this, we consider the deep quench protocol, wherein initially decoupled lattice and electronic subsystems (with $g_{i}=0$) are abruptly coupled through a large nonzero interaction $g_{f}\neq 0$. The ensuing dynamics are characterized by pronounced chaotic behavior. Owing to the accumulation of prediction errors, it is inherently infeasible for machine learning models to resolve long-time trajectories of such chaotic systems. Nonetheless, we demonstrate that a recently proposed Physics Aware Recurrent Convolution (PARC) architecture is capable of faithfully reproducing the statistical properties, or “climate,” of the spatiotemporal dynamics generated by the deep quench.
IV.1 PARC architecture
We begin by discussing the PARC neural network architecture, originally introduced to model the dynamics of highly nonlinear PDEs [17]. Rather than employing a single neural network to approximate the mapping $\mathbf{u}(t)\mapsto\mathbf{u}(t+\Delta t)$, PARC introduces two convolutional neural networks, which play the roles of a differentiator and an integrator, respectively, in analogy with the two stages of a finite-difference scheme for solving differential equations. This design is directly motivated by the numerical integration of differential equations over a finite time interval $\Delta t$.
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\int_{t}^{t+\Delta t}\frac{d\mathbf{u}}{dt}dt,$ (15)
[…]
In the shallow quench case discussed in Sec. III, a single CNN is introduced to approximate the second integral term in the above equation. The motivation of the PARC approach is to incorporate the two-step operations: differentiation followed by integration, into the recurrent model. To this end, we first introduce a CNN model $\boldsymbol{\mathcal{D}}[\cdot|\bm{\phi}]$ to approximate the differential operation
$\boldsymbol{\mathcal{D}}[\mathbf{u}|\bm{\phi}]\approx\frac{d\mathbf{u}}{dt},$ (16)
[…]
$\boldsymbol{\mathcal{S}}[\bm{f}|\bm{\theta}]\approx\int_{t}^{t+\Delta t}\bm{f}(t^{\prime})dt^{\prime},$ (17)
where $\bm{\theta}$ are the corresponding trainable parameters. The increment of the state vector over $\Delta t$ can then be expressed as the composition of differential and integral operators. Explicitly, the mapping over one time-step is given by
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\boldsymbol{\mathcal{S}}\Bigl[\boldsymbol{\mathcal{D}}\bigl[\mathbf{u}(t)\,\big|\,\bm{\phi}\bigr]\,\Big|\,\bm{\theta}\Bigr].$ (18)
Compared to the single-CNN recurrent structure in Eq. (10), this approach decomposes the single mapping $\boldsymbol{\mathcal{N}}$ into two distinct CNN models, corresponding to the differentiator and integrator operators, respectively.
Based on this physics motivated architecture, the PARC loss function extends the standard CNN loss Eq. (III.1) by adding a differentiator loss component. Let ${d\hat{\mathbf{w}}_{m}}/{dt}$ represent the mid-point (between the input time step and the prediction time step) derivative of the $m$-th component predicted by the CNN differentiator $\boldsymbol{\mathcal{D}}[\mathbf{u}(t)\,|\,\bm{\phi}]$, and let $\delta\hat{\mathbf{w}}_{m}$ be the update or difference of the $m$-th component predicted by the CNN integrator $\boldsymbol{\mathcal{S}}\bigl[\boldsymbol{\mathcal{D}}[\mathbf{u}(t)\,|\,\bm{\phi}]\,\big|\,\bm{\theta}\bigr]$. The PARC loss function becomes:
$\mathcal{L}(\bm{\phi},\bm{\theta}\,|\,\mathbf{u})=\sum_{m}\Biggl(\sum_{t}\norm{\frac{d\mathbf{w}_{m}}{dt}\bigg|_{t+\Delta t/2}-\frac{d\hat{\mathbf{w}}_{m}}{dt}\bigg|_{t+\Delta t/2}}_{2}$
$\qquad\qquad+\sum_{t}\norm{\mathbf{w}_{m}(t+\Delta t)-\mathbf{w}(t)-\delta\hat{\mathbf{w}}_{m}(t)}_{2}\Biggr),$
where $\mathbf{u}(t)=\{\mathbf{w}_{m}(t)\}$ denotes the ground-truth trajectory of the state vector obtained from direct numerical simulations.
Figure 6: Snapshots of phonon displacements $Q_{i}$, momentum $P_{i}$, and electron density matrix $\rho_{ij}$ at various times after a deep quench from $g_{i}=0.0$ to $g_{f}=1.0$ of the semi-classical Holstein model on a $L=16$ chain. The ground truth on the top row is obtained from the 4th-order Runge-Kutta integration of the governing dynamics Eqs. (3) and (6). The bottom row shows predictions from the trained PARC-based model. Only the real component of the complex-valued $\rho$ is shown. The steps are in units of prediction time steps, which for the deep quench case is set to $\Delta t=2.56$ time units. Step $25$ correlates to $64$ time units after the quench, which is the point by which the system enters its post-transient regime. The model predictions start at prediction step $25$ and continue for $40$ predictions steps until step $65$.
IV.2 Architecture Implementation Details
We next discuss details of the PARC implementation. A schematic diagram of the recurrent structure of our PARC-based model is shown in FIG. 5. As described earlier, PARC introduces two CNNs corresponding to the differentiator and integrator steps of a finite-difference method for solving differential equations.
A single PARC-based neural network is constructed to evolve all three dynamical variables, $\rho$, $Q$, and $P$. Both the differentiator and integrator components are structurally identical CNNs, sharing the same architecture as the standard CNN described in Sec. III.2. The input is prepared as follows: the complex-valued $L\times L$ matrix $\rho$ is decomposed into a real-valued $L\times L\times 2$ tensor, while $Q$ and $P$ are placed along the diagonals of $L\times L\times 1$ tensors. Concatenating these yields an $L\times L\times 4$ input tensor.
Each CNN (differentiator or integrator) maps this input to an output tensor of the same dimension. The physical quantities of interest - $\frac{d\rho}{dt}$, $\frac{dQ}{dt}$, $\frac{dP}{dt}$ in the case of the differentiator, and $\Delta\rho$, $\Delta Q$, $\Delta P$ in the case of the integrator - are then obtained by reversing the embedding procedure: recombining the first two channels into a complex $L\times L$ matrix for $\rho$ and extracting $Q$ and $P$ from the diagonals of the remaining channels. A schematic of this data flow is shown in Fig. 2. Additionally, we calculate data scaling coefficients from the data and integrate them into the model’s inputs and outputs as a form of data normalization. Although they are part of the model, we did not include them in Fig. 2 to reduce complexity. Further details regarding the data scaling coefficients can be found in appendix A.
IV.3 Deep Quench Data
As with the shallow quench data discussed in section III.3, the deep quench training dataset was obtained from direct numerical simulations of the semi-classical Holstein model. In the deep quench scenario, the coupling is initially turned off $g_{i}=0$, which means the corresponding ground state is free electron gas on a chain decoupled from a set of independent simple harmonic oscillators in equilibrium $Q_{i}=P_{i}=0$. Naturally, there is no CDW order in this initial state. The electron-phonon coupling is then suddenly switch on to $g_{f}>0$ at time $t=0$. The quench dynamics in this scenario is dominated by the emergence of CDW orders induced by the nonzero $g_{f}$. The onset of the CDW order is a stochastic process. First local CDW order is initiated through nucleation process at random seeds. This is then followed by the growth and merger of CDW domains. This scenario is akin to the phase-ordering process when a system is suddenly quenched from a high-temperature random state into a symmetry-breaking phase. As a result, the deep-quench process is very sensitive to random fluctuations in the initial states. For a controlled approach to generate dataset, $Q_{i}$ are sampled from a normal distribution of zero mean and a standard deviation of $10^{-4}$. Contrary to the shallow quench scenario discussed above, a wide variety of dataset with different system trajectories can be generated from the deep quench simulations.
To generate the deep-quench dataset, we performed a quench from $g_{i}=0.0$ to $g_{f}=1.0$, yielding a total of $1228$ trajectories. While the integration time step used to generate the data is still $\delta t=0.01$ time units, we set the prediction time step to $\Delta t=256\,\delta t=2.56$ time units. For the deep quench, snapshot capture began only after the first $64$ time units, by which point the system had reached its post-transient regime. After this, we capture a total of $1001$ prediction time steps (an initial state and $1000$ steps), totaling a time domain per trajectory of $1000\,\cdot\,2.56=2560$ time units. To enable the calculation of derivatives between prediction time steps - as seen in the differentiator term of the PARC loss function as shown in Eq. (IV.1) -, we additionally record the intermediate states between successive prediction time steps. Finally, in a similar fashion to the generation of the shallow quench data, we offset the starting point of each trajectory by a certain number of integration time steps to provide uniform temporal sampling throughout the prediction time interval of $\Delta t=2.56$.
IV.4 Training Details
The training procedure for the PARC-based model is a direct extension of that for the standard CNN model described in Sec. III.4, with only minor modifications to accommodate the additional structure of PARC.
Given datasets in the form of exact system trajectories $\mathbf{u}(t)=[\rho(t),Q(t),P(t)]$, the PARC loss function, Eq. (IV.1), is minimized via gradient descent to obtain the optimized parameters $\bm{\phi}^{*}$ and $\bm{\theta}^{*}$. Training proceeds with multi-step prediction: for each trajectory segment, $N$-step predictions are compared against the corresponding ground-truth states, while the differentiator simultaneously produces $N$ intermediate derivative estimates. These are evaluated against derivatives computed from the mid-interval states, providing the additional supervision required by the PARC framework.
Figure 7: Comparison of $\Delta_{\rho}$ traces between ground truth data and prediction trajectories starting from the same initial condition. The ground truth trajectories/initial conditions are randomly selected from a deep quench dataset with system size $32$. Blue traces are from ground truth data and red traces are from model predictions. Model predictions were made for $1000$ steps after the model was provided with a ground truth starting state. Steps are in units of $\Delta t=2.56$ time units. The model predictions adhere well to the ground truth traces in the short-term, but diverge in the long-term.
V Deep Quench Results
FIG. 6 visualizes a ground-truth trajectory and predictions from a PARC-based model trained on a deep quench dataset with system size $L=16$. The model was given the state at prediction step $25$ as input (which corresponds to the start of the saved deep quench dataset following the initial 64 time units of numerical integration during the system’s transient regime) and then predicted for $40$ steps thereafter to create the visualization. As previously noted, the deep quench trajectories are highly sensitive to initial condition, and small perturbations to the initial condition result in different system trajectories. All trajectories used in the analysis of the deep quench case belonged to a test set not shared with the training set. Additionally, both ground truth trajectories and predictions have initial states starting after the transient regime, which, as noted in Sec. IV.3, begins $64$ time units after the quench.
The shared initial state in both cases is characterized by nearly constant diagonal elements $\rho_{ii}\sim\text{const.}$, indicating a uniform charge distribution, together with approximately uniform lattice displacements $Q$ and momenta $P$. As discussed above, this configuration corresponds to decoupled electron and lattice subsystems close to their respective ground states. Once the coupling is switched on at $g_{f}=1$, energetic considerations suggest that the system tends to develop CDW order. Because the formation of CDW order in distant regions is largely independent, owing to the local nature of the symmetry-breaking process, the system is expected to evolve into an inhomogeneous state consisting of multiple CDW domains of opposite sign, separated by domain walls. Indeed, while the electron density remains relatively uniform, as indicated by the diagonal line of the density matrix, a pronounced inhomogeneous CDW state emerges at later times following the quench (see Fig. 6), becoming visible by time-step $25\Delta t$.
Remarkably, even for such strongly inhomogeneous space-time dynamics, the ML predictions remain in close agreement with the ground truth at $40\,\Delta t$ after the beginning the prediction. This highlights the ability of the model to capture highly nontrivial dynamical features well beyond the initial stages of evolution. Nevertheless, as with any statistical learning approach, prediction errors are unavoidable. While such errors may remain small at short and intermediate times, their gradual accumulation is inevitable and eventually compromises the reliability of long-time forecasts. This limitation is clearly illustrated in Fig. 7, which compares the time traces of the CDW order parameter $\Delta_{\rho}(t)$ from four independent exact simulations with those obtained from the PARC model. Importantly, this issue is not specific to the present implementation but reflects a fundamental challenge of ML-based dynamical modeling, where small stepwise inaccuracies compound over time and obscure the true asymptotic behavior of the system.
Long-term prediction is particularly difficult in chaotic dynamical systems due to their extreme sensitivity to small variations in initial conditions. The post-quench dynamics of the semi-classical Holstein model exhibits such chaotic behavior, as shown in Fig. 7: four independent $g_{i}=0\to g_{f}=1$ quench simulations initiated with slightly different stochastic noise profiles produce markedly distinct trajectories of $\Delta_{\rho}(t)$ (blue curves). Beyond this hypersensitivity to initial conditions, the CDW order parameter exhibits no emergent structure or reproducible dynamical pattern. As a result, even a highly optimized model will begin to diverge from a ground truth trajectory over many steps.
Figure 8: Ground truth vs predicted autocorrelation trajectories of both $\Delta_{\rho}$ and $\Delta_{Q}$. Values of $\tau$ up to a maximum of $250$ were used out of $1000$ total steps in the trajectories. The same starting states from the ground truth trajectories were used for the predicted trajectories. There were $256$ each of ground truth and predicted trajectories. The system size is $32$.
On the other hand, while the ML predictions drift away from the ground-truth trajectories, the predicted time traces seem to retain the same statistical characteristics, implying that our ML models could successfully capture the underlying chaotic attractor dynamics. To demonstrate this, we consider the autocorrelation function for the order parameters ($X=\Delta_{\rho}$ or $\Delta_{Q}$):
$A(\tau)=\frac{\langle X(\tau_{0})X(\tau_{0}+\tau)\rangle-\langle X(\tau_{0})\rangle}{\langle X(\tau_{0})^{2}\rangle-\langle X(\tau_{0})\rangle}^{2},$ (19)
Figure 9: An illustration showing the application of data scaling coefficients on the input and output of the model.
where $\tau$ is the time lag, and the brackets $\langle\cdots\rangle$ denote averaging over both the initial time $t_{0}$ and an ensemble of independent runs, which together approximate an average over the invariant measure of the time series. The autocorrelation function provides a quantitative measure of temporal correlations in a dynamical trajectory and is particularly useful for diagnosing chaotic behavior. By definition, $A(0)=1$, since a trajectory is perfectly correlated with itself at zero lag. As $\tau$ increases, $A(\tau)$ decays, reflecting the gradual loss of temporal memory. For trajectories evolving on a chaotic attractor, $A(\tau)$ typically approaches zero at large lag times, consistent with the system’s sensitivity to initial conditions and its ergodic exploration of the attractor. In this sense, the autocorrelation function probes not only local dynamical behavior but also the invariant measure that governs long-time statistics.
We apply this analysis to the charge-density-wave order parameter $\Delta_{\rho}(t)$ and the lattice displacement $\Delta_{Q}(t)$, comparing ground-truth trajectories with those generated by the ML model. For each case, $256$ independent trajectories of length $1000$ time steps were analyzed, with lag values considered up to $250$ steps. The resulting autocorrelation curves are shown in FIG. 8. Strikingly, the predicted trajectories reproduce the autocorrelation functions of the ground truth with high fidelity. This agreement demonstrates that the ML model has not only learned to reproduce individual short-time trajectories, but also captured the ergodic properties and invariant statistical structure of the underlying chaotic dynamics.
Figure 10: The prediction error of different models over a time interval of $81.92$ time units. Different models were trained with different prediction intervals ranging from $\Delta=0.32$ to $\Delta=10.24$. Models with shorter prediction intervals required more steps to predict over the same interval of time. The ”Total Avg RMSE” represents the sum of the average RMSE for each of the components $\rho$, $Q$, and $P$ over $256K$ random prediction targets from a test set.
VI Summary and Outlook
We have demonstrated that recurrent convolutional architectures can effectively model the nonequilibrium dynamics of the one-dimensional semi-classical Holstein model across distinct quench protocols. In the case of the shallow quench, where the dynamics are largely deterministic and exhibit simple periodic structure, we showed that even an extremely compact recurrent CNN, with only a few thousand trainable parameters, is sufficient to accurately reproduce the system’s time evolution. This highlights the efficiency of standard deep-learning architectures in capturing relatively simple dynamical regimes.
By contrast, the deep quench scenario presents a far greater challenge, as the sudden onset of strong electron-phonon coupling drives chaotic spatiotemporal dynamics that defy accurate trajectory-level prediction. For this regime, we introduced a Physics Aware Recurrent Convolutional (PARC) architecture, trained solely on sequential state data, and demonstrated its ability to faithfully reproduce the long-term statistical “climate” of the dynamics. The efficacy of the PARC model derives from its differentiator-integrator structure, which encodes a physics-aware inductive bias directly into the network. This hybrid design augments the CNN’s capacity to learn local hierarchical correlations with a structural prior aligned to the underlying equations of motion.
[…]